Media services charge $40 for a phone and $30/month for its economy pla. The equation c=30t+40 describes the total cost, c, of operating a media services phone for t months. The total cost for 7 months of service$________, and If a cutomer has only $110 available, how many months of service can she receive?
Answer by jorel1380(3719)

40+30(7)=210+40=$250.
110-40=70/30=2.33--She can only receive 2 1/3 months for $110..